OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 4x − 2y if x = −1 and y = −5

OpenStudy (mathstudent55):

Just plug in -1 for x and -5 for y. 4x - 2y = 4 * (-1) - 2 * (-5) Now do the two multiplications first, then the subtraction.

OpenStudy (anonymous):

this is confusing

OpenStudy (mathstudent55):

What is 4 times negative 1?

OpenStudy (anonymous):

-4

OpenStudy (mathstudent55):

good What is 2 times -5?

OpenStudy (anonymous):

-10

OpenStudy (mathstudent55):

good

OpenStudy (mathstudent55):

That means that 4 * (-1) - 2 * (-5) is the same as -4 - (-10) How do you do that subtraction?

OpenStudy (mathstudent55):

Rule of subtraction: To subtract a number from another number, change the subtraction into addition of the opposite. For example, 5 - (-6) = 5 minus the opposite of -6, or 6, so 5 - (-6) = 5 + 6

OpenStudy (anonymous):

so its -14

OpenStudy (mathstudent55):

-4 - (-10) turns into addition of the opposite: -4 + 10

OpenStudy (mathstudent55):

6

OpenStudy (anonymous):

6?